OpenStudy (z4k4r1y4):

Solve for x: 2cosh(x)=0

OpenStudy (z4k4r1y4):

cosh(x)=1/2(e^x+e^-x)

OpenStudy (z4k4r1y4):

so 2coshx= e^x+e^-x which = 0

OpenStudy (mathmate):

and \(e^x >0~~\forall x\in R\) so there is no real root.

OpenStudy (z4k4r1y4):

exactly the answer I got!

OpenStudy (z4k4r1y4):

thank you for verifying.